i cant figure out how to solve this radical equation square root of 20-8x=x
Answer by stanbon(57282)

square root of 20-8x=x
---
If you meant sqrt(20-8x) = x
Square both sides to get: 20-8x = x^2
x^2+8x-20 = 0
(x+10)(x-2) = 0
Only solution:
x = 2 because sqrt(20-8x) must not be negative.
==========================
Cheers,
Stan H.